RYDBERG STATES OF CALCIUM MONOCHLORIDE

\underline{JASON O. CLEVENGER}, XING JIANG, SERGEY I. PANOV, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139^{a}.

Experimental results from our continuing study of the Rydberg states of Calcium Monochloride will be presented. Previously we have characterized the quantum defects and predissociation mechanisms of core-penetrating ^{2}Σ^{+} states of CaCl in the n* = 3-7 region.^{a}^{,} ^{b} Questions unanswered by this work propelled further studies of the low-n* region (n* = 3-5) by REMPI and ion-dip techniques, as this region is expected to be extensively predissociated by ^{2}Σ^{+} and ^{2}Π repulsive states. Interest in high-n* (n* > 15) Rydberg states of CaCl has also fueled further REMPI experiments near the first ionization threshold (48 491 cm^{−1}). Studies of the low-n* region failed to reveal several predicted members of known core-penetrating ^{2}Σ^{+} Rydberg series (n* = 0.16, 0.49, 0.68 ^{2}Σ^{+} and 0.30 ^{2}Π), while preliminary data for the high-n* region above the v^{+} = 0 IP indicate the presence of vibrationally autoionizing Rydberg states converging to v^{+} > 0 vibrational levels of the CaCl^{+} ion X ^{1}Σ^{+} state. Analysis is ongoing.

aThis research is supported by NSF Grant CHE97-30852.

aJ. Li, Y. Liu, D. B. Moss, C. M. Gittins, N. A. Gittins, and R. W. Field, J. Mol. Spec. \underline{193}, 403 (1999).

bJ. O. Clevenger, N. A. Harris, R. W. Field, J. Li, J. Mol. Spec. \underline{193}, 412 (1999).